# A recursive approach for the determination of the nice sections of width three which have a 4-crown stack as retract Version 8-3

### v2

@@ Abstract
The characterization of the finite minimal automorphic posets of width three is still an open problem. Niederle has shown that this task can be reduced to the characterization of the nice sections of width three having a non-trivial tower of nice sections as retract. In our article, we develop a recursive approach for the determination of those of them which have a 4-crown stack as retract. We apply the approach on a sub-class $\mathfrak{N}_{2}$ of nice sections of width three and determine all posets in $\mathfrak{N}_{2}$ with height up to six having a 4-crown stack as retract. For each integer $n\geq 2$, the class $\mathfrak{N}_{2}$ contains $2^{n-2}$ different isomorphism types of posets of height $n$.
@@ 1 Introduction
However, the description or characterization of minimal automorphic posets is still an open field. (Unless otherwise stated, “poset” always means “finite poset” in this article.) Since the pioneering work of Rival [7, 8], we know that a poset of height one does not have the fixed point property iff it contains a crown, because a crown of minimal length will be a retract of it. Brualdi and Dias Da Silva [1] generalized this result by showing that a poset does not have the fixed point property iff it has a “generalized crown” as retract. With respect to posets of small width, Fofanova and Rutkowski [4] showed that a poset of width two does not have the fixed point property iff it has a 4-crown stack as retract. (The result extends even to chain-complete infinite posets.) Niederle [5] characterized in 1989 posets of width three not having the fixed point property by having a non-trivial retract being a “tower of nice sections”. Later on, in 2008, he extended his result to posets of width four [6].
@@
Due to the results of Niederle [5], the study of the minimal automorphic posets of width three can be restricted to the investigation of the nice sections of width three. In our article, we develop and apply a recursive approach for their determination. After the preparatory Section 2, we deal with the classes $\mathfrak{N}$ of nice sections and $\mathfrak{C}$ of crowned sections in Section 3; $\mathfrak{N}$ is a proper subclass of $\mathfrak{C}$. We see that a poset $P\in\mathfrak{C}$ has a 4-crown stack as retract iff it has such a retract containing a split level, and we prove several results about the existence of such levels. We see in Lemma 3.11 that a split level of $R$ corresponds in $P$ to a certain decomposition by a retractive split, and based on this observation we develop at the end of Section 3 a recursive approach for the determination of the nice sections having a 4-crown stack as retract.
In order to test the approach, we introduce the class $\mathfrak{N}_{2}$ of nice sections with horizon two in Section 4. For each integer $n\geq 2$, this class contains $2^{n-2}$ isomorphism types of posets of height $n$; the isomorphism types are described in Theorem 4.2. We characterize in Theorem 4.5 the retractive splits of posets $P\in\mathfrak{N}_{2}$ yielding a 4-crown stack as retract. Using this result, we determine in Section 5 by means of the recursive approach all posets $P\in\mathfrak{N}_{2}$ of height up to six which have a 4-crown stack as retract.
The characterization of the minimal automorphic posets of width three is regarded as a difficult task [9, p. 98], in parts even as “intractable” [3, p. 126]. Also the sparse sequence of publications indicates these difficulties. Nevertheless, it gives some confidence that the nice section of width three used by Schröder [9, p. 99] for the illustration of the problems and considered as “nasty” by Farley [3, p. 136] is handled without a problem by the machinery developed in this article. It is the poset coded as 1011 in Section 5 and shown in Figure 9.
@@ 2 Preparation
In this section, we introduce our notation and recapitulate definitions of structures being in the focus of our investigation. For all other terms of order theory, the reader is referred to standard textbooks as [9]. For integers $i\leq j$, we write $[i,j]$ for the interval $\{i,\ldots,j\}$.
@@
Given disjoint posets $P=(X,\leq_{P})$ and $Q=(Y,\leq_{Q})$, their ordinal sum is the poset $P\oplus Q:=(X\cup Y,\leq_{P}\cup\leq_{Q}\cup\;(X\times Y))$. For $P$ and $Q$ being disjoint antichains of size 2 or 3, we simply write “#P#Q\#P\#Q” as shortcut for the isomorphism type of $P\oplus Q$, e.g., $22$ indicates a 4-crown. Additionally, we write $2C$ and $3C$ for a poset consisting of two (three) disjoint chains of height 1 without any additional comparabilities between different points. The shortcuts are illustrated in Figure 1.
@@
An order homomorphism $r:P\rightarrow P$ is called a retraction of the poset $P$ if $r$ is idempotent, and an induced sub-poset $R$ of $P$ is called a retract of $P$ if a retraction $r:P\rightarrow P$ exists with $r[X]=R$. For the sake of simplicity, we identify $r$ with its post-restriction and write $r:P\rightarrow R$. A poset $P$ has the fixed point property iff every retract of $P$ has the fixed point property [9, Th. 4.8]. We need the following well-known result about the existence of spanning retracts:
@@
Let $P$ be be a finite poset and $r:P\rightarrow R$ a retraction. There exists a retraction $v:P\rightarrow V$ with $V\simeq R$, $\min V\subseteq\min P$, $\max V\subseteq\max P$, and $V(1\rightarrow h_{V}-1)=R(1\rightarrow h_{R}-1)$.
@@
3 Nice sections
3.1 Definition and basic properties
@@
Niederle [5, p. 121 and p. 125] has shown that the decomposition of a tower of sections into sections is unique, and that a poset of width at most three has not the fixed point property iff it has a proper retract being a tower of nice sections. We conclude that an automorphic poset $P$ of width at most 3 not being a tower of nice sections is not minimal automorphic. Moreover, if $P=S_{1}\oplus\cdots\oplus S_{n}$ is a tower of nice sections $S_{1},\ldots S_{n}$, it is not hard to see that $P$ is minimal automorphic iff $S_{i}$ is minimal automorphic for every $i\in[1,n]$. The objects we have to deal with in investigating minimal automorphic posets of width three are thus the nice sections of width three.
@@ 3.2 Crowned sections and split levels
Our goal is a recursive approach for the determination of the nice sections of width three having a 4-crown stack as retract. However, our results will concern a larger class of sections containing the nice sections of width three as a proper subclass:
@@
We agree on that for the rest of this section the poset $P$ is a crowned section and $r:P\rightarrow R$ is a retraction onto a 4-crown stack. This implies $h_{R}\geq 2$: Suppose to the contrary that $R$ is a 4-crown. For the corresponding spanning retraction $r^{\prime}:P\rightarrow R^{\prime}$, the poset $Q:=P(0,1)\cup P(h_{P})$ has the retraction $r^{\prime}|_{Q}:Q\rightarrow R^{\prime}$ onto the 4-crown $R^{\prime}$. But this is impossible because $Q$ is a 6-crown or a 6-crown stack of height 2, or it contains a point $x\in Q(1)$ with $R^{\prime}(0)<x<R^{\prime}(1)$.
In our investigation of split levels, the level sets $R(0)$, $R(1)$, $R(h_{R}-1)$, and $R(h_{R})$ frequently will play a special role. We notate the results for $R(0)$ and $R(1)$ only; the results for $R(h_{R})$ and $R(h_{R}-1)$ are dual.
@@
Lemma 2.1 and Lemma 3.4 together yield that a crowned section has a 4-crown stack as retract iff it has a 4-crown stack as retract containing a split level. On the one hand, this inconspicuous result brings our recursive approach into motion in Section 3.3. On the other hand, this approach becomes more effective and the concept of split levels more applicable if we can show that frequently there exist additional split levels not induced by the action of $r$ on $P(0)$ and $P(h_{P})$. In order to separate induced split levels from independent ones, we discriminate two cases:
  • $R(0)\subset P(0)$: Looking at the proof Lemma 3.4, we realize that if $R(0)$ is not a down-split level, then $R(1)$ is forced to be an up-split level.
  • $R(0)\not\subset P(0)$: Due to (1), there exists a point $a\in R(0)\cap P(0)$. For the point $b\in R(0)\setminus\{a\}$, we have $r[P(0)\setminus\{a\}]=\{b\}$, hence $r[P(1)\setminus\{c_{1,\gamma(b)}\}]\subseteq R(1\rightarrow h_{R})$. Nevertheless, two alternatives remain for the level set $R(1)$: it can be an up-split level or no split level at all.
In consequence, we define
Definition 3.5.
We call a split level $R(\ell)$ induced if
  • $\ell=0$ or $\ell=h_{R}$;
  • $\ell=1$, $R(0)\subset P(0)$, and $r[P(0)]\not\subset R(0)$;
  • $\ell=h_{R}-1$, $R(h_{R})\subset P(h_{P})$, and $r[P(h_{P})]\not\subset R(h_{R})$.
Otherwise, we call $R(\ell)$ independent.
For the rest of this section, we investigate structures which ensure the existence of independent split levels. We start with
@@
    $\ell=0$: The level set $R(1)$ is an independent up-split level if
@@
1) We have $\lambda(a)=0$ according to (1). Due to $P(0)\setminus\{a\}<b$, we have $r(x)=b$ for all $x\in P(0)\setminus\{a\}$, thus $R(0)<r(y)$ for all $y\in P(\min\lambda[R(1)])$, and $R(1)$ is an up-split level. It is an independent one because of $R(0)\not\subset P(0)$.
@@
2) If $P(\lambda(b),\lambda(v))\simeq 33$, then $R(\ell)$ is a down-split level, $R(\ell+1)$ is an up-split level, and at least one of these split levels is independent.
@@
1) The points $a,c_{\lambda(a),\gamma(b)}$ and $w,c_{\lambda(w),\gamma(v)}$ form a 4-crown.
2) The assumption yields $P(\lambda(b))<R(\ell+1)$ and $R(\ell)<P(\lambda(v))$. In consequence, $R(\ell)$ is a down-split level and $R(\ell+1)$ an up-split level. For $1\leq\ell\leq h_{R}-2$, both split levels are independent, and for $\ell=0$, $R(0)$ being a down-split level makes $R(1)$ independent. For $\ell=h_{R}-1$, apply the dual argument.
3) The point $v$ has two lower covers $b$ and $z$ in $P(\lambda(b)),\lambda(v))$, and we have $z<R(\ell+1)$, hence $r(z)\in R(0\rightarrow\ell)$. For the point $y\in P(\lambda(b))\setminus\{b,z\}$, we have $y<P(\lambda((v))\setminus\{v\}$. If $r(y)\in R(0\rightarrow\ell)$, then $R(\ell)$ is a down-split level, and in the case of $r(y)\in R(\ell+1\rightarrow h_{R})$, the level $R(\ell+1)$ is an up-split level. The addendum follows directly from the definition of an independent split level.
@@
The relation $\lambda(a)<\lambda(b)$ is due to Lemma 3.4. Because of the dual of Lemma 3.7.3, the level pair $P(\lambda(a),\lambda(v))$ cannot be of type 33, and $P(\lambda(a),\lambda(v))$ must be a 6-crown. This implies $P(\lambda(a),\lambda(b))\simeq 3C$, and $a<v$ yields $a<b^{\prime}$. Due to Lemma 3.7.1, the type of $P(\lambda(a),\lambda(w)))$ is 33. Finally, $r(v^{\prime})\not=b$ would yield $r(v^{\prime})\in R(\ell+1\rightarrow h_{R})$, and depending on the image of the point in $P(\lambda(b))\setminus\{b,b^{\prime}\}$, the level set $R(\ell)$ would be a down-split level or $R(\ell+1)$ an up-split level. The rest is dual.
@@
the level set $R(\ell)$ is a down-split level or the level set $R(\ell+1)$ is an up-split level, and at least one of these split levels is independent.
@@
Suppose thus that $P(\lambda(b),\lambda(v))$ is a 6-crown. If the type of $P(\lambda(a),\lambda(v))$ or $P(\lambda(b),\lambda(w))$ is 33, then apply Lemma 3.7.3 or its dual. Otherwise, we must have $\lambda(v)=k+1=\lambda(w)$. According to Lemma 3.4, $R(\ell+1)$ is an up-split level, hence independent, or a down-split level, which is independent for $\ell+2<h_{R}$.
Let thus $R(\ell+1)$ be a down-split level and $\ell+2=h_{R}$. Due to $R(\ell+1)\subset P(k+1)$, we have $R(h_{R})\not\subset P(k+2)$, thus $h_{P}\geq k+3$, and $R(\ell+1)<P(h_{P})$ follows. We conclude $r[P(h_{R})]=R(h_{R})$, and $R(\ell+1)$ is independent.
@@ 3.3 A recursive approach
This result gives raise to a recursive approach to determine the crowned sections having a 4-crown stack as retract. Let $\mathfrak{S}$ be a subclass of $\mathfrak{C}$ closed under duality, e.g., the subclass of nice sections of width three. For a retractive split of a poset $P\in\mathfrak{S}$, we need a $t$-base $V$ and an $s$-base $W$ of a poset $P\in\mathfrak{S}$ of width 3, i.e.,
  • $V$ is a lower segment or an upper segment of $P$ having a retraction $t$ onto a 2-antichain or a 4-crown stack $T$.
  • $W$ is an upper segment of $P$ having a retraction of $W\setminus D$ onto a 2-antichain or a 4-crown stack $S$ for a down-set $D\subset W$, or $W$ is a lower segment of $P$ having a retraction of $W\setminus U$ onto a 2-antichain or a 4-crown stack $S$ for an up-set $U\subset W$.
@@
In order to test the approach we need a sub-class of $\mathfrak{C}$ combining plenty of interesting posets with good structural access. We choose the class of the nice sections with horizon two. We introduce them in Section 4, and in Section 5 we identify all of them with height up to six having a 4-crown stack as retract.

### v3

@@ Abstract
The characterization of the finite minimal automorphic posets of width three is still an open problem. Niederle has shown that this task can be reduced to the characterization of the nice sections of width three which have a non-trivial tower of nice sections as retract. In our article, we develop a recursive approach for the determination of those nice sections of width three which have a 4-crown stack as retract. We apply the approach on a sub-class $\mathfrak{N}_{2}$ of nice sections of width three and determine all posets in $\mathfrak{N}_{2}$ with height up to six which have a 4-crown stack as retract. For each integer $n\geq 2$, the class $\mathfrak{N}_{2}$ contains $2^{n-2}$ different isomorphism types of posets of height $n$.
@@ 1 Introduction
However, the description or characterization of minimal automorphic posets is still an open field. (Unless otherwise stated, “poset” always means “finite poset” in this article.) Since the pioneering work of Rival [7, 8], we know that a poset of height one does not have the fixed point property iff it contains a crown, because a crown of minimal length will be a retract of it. Brualdi and Dias Da Silva [1] generalized this result by showing that a poset does not have the fixed point property iff it has a “generalized crown” as retract. With respect to posets of small width, Fofanova and Rutkowski [4] showed that a poset of width two does not have the fixed point property iff it has a 4-crown stack as retract. (The result extends even to chain-complete infinite posets.) Niederle [5] proved in 1989 that a poset of width three does not have the fixed point property iff it has a “tower of nice sections” as non-trivial retract. Later on, in 2008, he extended his result to posets of width four [6].
@@
Due to the results of Niederle [5], the study of the minimal automorphic posets of width three can be restricted to the investigation of the nice sections of width three. In our article, we develop and apply a recursive approach for the determination of those nice sections of width three which have a 4-crown stack as retract. After the preparatory Section 2, we deal with the classes $\mathfrak{N}$ of nice sections and $\mathfrak{C}$ of crowned sections in Section 3; $\mathfrak{N}$ is a proper subclass of $\mathfrak{C}$. We see that a poset $P\in\mathfrak{C}$ has a 4-crown stack as retract iff it has such a retract which contains a split level, and we prove several results about the existence of such levels. We see in Lemma 3.10 that a split level of $R$ corresponds in $P$ to a certain decomposition by a retractive split, and based on this observation we develop at the end of Section 3 a recursive approach for the determination of the nice sections with a 4-crown stack as retract.
In order to test the approach, we introduce the class $\mathfrak{N}_{2}$ of nice sections with horizon two in Section 4. For each integer $n\geq 2$, this class contains $2^{n-2}$ isomorphism types of posets of height $n$; the isomorphism types are described in Theorem 4.2 which generalizes a result of Farley [3, Prop. 4.1]. We characterize in Theorem 4.5 the retractive splits of posets $P\in\mathfrak{N}_{2}$ which yield a 4-crown stack as retract. Using this result, we determine in Section 5 by means of the recursive approach all posets $P\in\mathfrak{N}_{2}$ of height up to six with a 4-crown stack as retract.
The characterization of the minimal automorphic posets of width three is regarded as a difficult task [9, p. 98], in parts even as “intractable” [3, p. 126]. Also the sparse sequence of publications indicates difficulties. Nevertheless, it gives some confidence that the nice section of width three used by Schröder [9, p. 99] for the illustration of the problems and considered as “nasty” by Farley [3, p. 136] is handled without a problem by the machinery developed in this article. It is the poset coded as 1011 in Section 5 and shown in Figure 9.
@@ 2 Preparation
In this section, we introduce our notation and recapitulate definitions of structures which are in the focus of our investigation. For all other terms of order theory, the reader is referred to standard textbooks as [9]. For integers $i\leq j$, we write $[i,j]$ for the interval $\{i,\ldots,j\}$.
@@
Given disjoint posets $P=(X,\leq_{P})$ and $Q=(Y,\leq_{Q})$, their ordinal sum is the poset $P\oplus Q:=(X\cup Y,\leq_{P}\cup\leq_{Q}\cup\;(X\times Y))$. For $P$ and $Q$ being disjoint antichains of size 2 or 3, we simply write “#P#Q\#P\#Q” as shortcut for the isomorphism type of $P\oplus Q$, e.g., $22$ indicates a 4-crown. Additionally, we write $2C$ and $3C$ for a poset which consists of two (three) disjoint chains of height 1 without any additional comparabilities between different points. The shortcuts are illustrated in Figure 1.
@@
An order homomorphism $r:P\rightarrow P$ is called a retraction of the poset $P$ if $r$ is idempotent, and an induced sub-poset $R$ of $P$ is called a retract of $P$ if a retraction $r:P\rightarrow P$ exists with $r[X]=R$. For the sake of simplicity, we always identify $r$ with its post-restriction and write $r:P\rightarrow R$. A poset $P$ has the fixed point property iff every retract of $P$ has the fixed point property [9, Th. 4.8]. We need the following well-known result about the existence of spanning retracts:
@@
Let $P$ be be a finite poset and $r:P\rightarrow R$ a retraction. There exists a retraction $v:P\rightarrow V$ with $V\simeq R$, $\min V\subseteq\min P$, and $\max V\subseteq\max P$.
@@
3 Nice sections and crowned sections
3.1 Nice sections
@@
Niederle [5, p. 121 and p. 125] has shown that the decomposition of a tower of sections into sections is unique, and that a poset of width at most three has not the fixed point property iff it has a proper retract which is a tower of nice sections. We conclude that a minimal automorphic poset $P$ of width at most three must be a tower of nice sections. Moreover, if $P=S_{1}\oplus\cdots\oplus S_{n}$ is a tower of nice sections $S_{1},\ldots S_{n}$, it is not hard to see that $P$ is minimal automorphic iff $S_{i}$ is minimal automorphic for every $i\in[1,n]$. The objects we have to deal with in investigating minimal automorphic posets of width three are thus the nice sections of width three.
@@ 3.2 Crowned sections and split levels
Our goal is a recursive approach for the determination of the nice sections of width three which have a 4-crown stack as retract. However, our results will concern a larger class of sections which contains the nice sections of width three as a proper subclass:
@@
We agree on that for the rest of this section the poset $P$ is a crowned section and $r:P\rightarrow R$ is a retraction onto a 4-crown stack. This implies $h_{R}\geq 2$. Suppose to the contrary that $R$ is a 4-crown. For the corresponding spanning retraction $r^{\prime}:P\rightarrow R^{\prime}$, the poset $Q:=P(0,1)\cup P(h_{P})$ has the retraction $r^{\prime}|_{Q}:Q\rightarrow R^{\prime}$ onto the 4-crown $R^{\prime}$. But this is impossible because $Q$ is a 6-crown or a 6-crown stack of height 2, or it contains a point $x\in Q(1)$ with $R^{\prime}(0)<x<R^{\prime}(1)$.
In our investigation of split levels, the level sets $R(0)$, $R(1)$, $R(h_{R}-1)$, and $R(h_{R})$ will frequently play a special role. We notate the results for $R(0)$ and $R(1)$ only; the results for $R(h_{R})$ and $R(h_{R}-1)$ are dual.
@@
Lemma 2.1 and Lemma 3.4 together yield that a crowned section has a 4-crown stack as retract iff it has a 4-crown stack as retract which contains a split level. On the one hand, this inconspicuous result brings our recursive approach into motion in Section 3.3. On the other hand, this approach becomes more effective and the concept of split levels more applicable if we can show that in most cases we can expect additional split levels which are not induced by the action of $r$ on $P(0)$ and $P(h_{P})$. We start with
@@
    $\ell=0$: The level set $R(1)$ is an up-split level if
@@
1) We have $\lambda(a)=0$ according to (1). Due to $P(0)\setminus\{a\}<b$, we have $r(x)=b$ for all $x\in P(0)\setminus\{a\}$, thus $R(0)<r(y)$ for all $y\in P(\min\lambda[R(1)])$, and $R(1)$ is an up-split level.
@@
2) If $P(\lambda(b),\lambda(v))\simeq 33$, then $R(\ell)$ is a down-split level and $R(\ell+1)$ is an up-split level..
@@
1) The points $a,c_{\lambda(a),\gamma(b)}$ and $w,c_{\lambda(w),\gamma(v)}$ form a 4-crown in $P(\lambda(a),\lambda(w))$.
2) The assumption yields $P(\lambda(b))<R(\ell+1)$ and $R(\ell)<P(\lambda(v))$.
3) The point $v$ has two lower covers $b$ and $z$ in $P(\lambda(b)),\lambda(v))$, and we have $z<R(\ell+1)$, hence $r(z)\in R(0\rightarrow\ell)$. For the point $y\in P(\lambda(b))\setminus\{b,z\}$, we have $y<P(\lambda((v))\setminus\{v\}$. If $r(y)\in R(0\rightarrow\ell)$, then $R(\ell)$ is a down-split level, and in the case of $r(y)\in R(\ell+1\rightarrow h_{R})$, the level $R(\ell+1)$ is an up-split level.
@@
The relation $\lambda(a)<\lambda(b)$ is due to Lemma 3.4. Because of the dual of Lemma 3.6.3, the level pair $P(\lambda(a),\lambda(v))$ cannot be of type 33, and $P(\lambda(a),\lambda(v))$ must be a 6-crown. This implies $P(\lambda(a),\lambda(b))\simeq 3C$, and $a<v$ yields $a<b^{\prime}$. Due to Lemma 3.6.1, the type of $P(\lambda(a),\lambda(w)))$ is 33. Finally, $r(v^{\prime})\not=b$ would yield $r(v^{\prime})\in R(\ell+1\rightarrow h_{R})$, and depending on the image of the point in $P(\lambda(v))\setminus\{v,v^{\prime}\}$, the level set $R(\ell)$ would be a down-split level or $R(\ell+1)$ an up-split level. The rest is dual.
@@
the level set $R(\ell)$ is a down-split level or the level set $R(\ell+1)$ is an up-split level.
@@
Suppose thus that $P(\lambda(b),\lambda(v))$ is a 6-crown. If the type of $P(\lambda(a),\lambda(v))$ or $P(\lambda(b),\lambda(w))$ is 33, then apply Lemma 3.6.3 or its dual. Otherwise, we must have $\lambda(v)=k+1=\lambda(w)$ and Lemma 3.4 delivers the result.
@@ 3.3 A recursive approach
This result gives raise to a recursive approach to determine the crowned sections which have a 4-crown stack as retract. Let $\mathfrak{S}$ be a subclass of $\mathfrak{C}$ closed under duality, e.g., the subclass of nice sections of width three. For a retractive split of a poset $P\in\mathfrak{S}$, we need a $t$-base $V$ and an $s$-base $W$ of a poset $P\in\mathfrak{S}$ of width 3, i.e.,
  • $V$ is a lower segment or an upper segment of $P$ which has a retraction $t$ onto a 2-antichain or a 4-crown stack $T$.
  • $W$ is an upper segment of $P$ which has a retraction of $W\setminus D$ onto a 2-antichain or a 4-crown stack $S$ for a down-set $D\subset W$, or $W$ is a lower segment of $P$ which has a retraction of $W\setminus U$ onto a 2-antichain or a 4-crown stack $S$ for an up-set $U\subset W$.
@@
In order to test the approach we need a sub-class of $\mathfrak{C}$ which combines plenty of interesting posets with good structural access. We choose the class of the nice sections with horizon two. We introduce them in the next section, and in Section 5, we identify all of them with height up to six with a 4-crown stack as retract.